# Breaking the News: A LLM-based Game where Players Act as Influencer or Debunker for Raising Awareness About Misinformation

### v1

@@ 3.4.1. Multiplayer Setup
Figure 6. Interaction System. Player 1 creates the online game room as the host and Player 2 joins. After Player 1 sends a request to GPT-4o API
, Player 2 can view it. Both players in the online room receive the API response, which is processed and updated on their own game screens. Then Player 2 takes their turn to input information, following the same process as Player 1. At the end of the round, all in-game events (including player inputs, API responses, time, etc.) are logged locally.
@@ 4.1.1. Questionnaires
Up to date, there is no standardized way of assessing the efficiency of prebunking interventions. Therefore, we specify the quantitative assessment based on Tifferet’s taxonomy of existing approaches (Tifferet, 2021) for evaluating misinformation susceptibility. Tifferet categorizes measurements into three main groups: performance tasks (how well users can discriminate between fake and real pieces of news), general media literacy assessment (how much a person knows about different aspects of misinformation), and behavioral assessment (how much a person would like to use different strategies to verify online information) (Tifferet, 2021). As Tifferet argues, these three aspects are complementary to understanding user’s susceptibility to misinformation and, therefore, evaluation of improvement in each of them could provide a full picture of our game’s efficiency. As the game was designed to present complex and multifaceted scenarios to the users, we wanted to assess how confident the participants were in their abilities to recognize misinformation before and after the intervention to see if the received knowledge increased or decreased their perceived self-efficacy in the topic. In the study, we used the validated questionnaires dedicated to these aspects.
Performance Assessment: Misinformation Susceptibility Test (MIST - 20). To assess changes in veracity discernment (the ability to recognize fake news from real news), we used the Misinformation Susceptibility Test (MIST) (Maertens et al., 2024). To date, the MIST inventory is the only fully validated misinformation susceptibility instrument. It takes into account the ability to recognize real news and fake news presented in equal proportion. The MIST framework is designed to allow for the comparison of results across different studies and interventions. The test has been implemented in multiple misinformation intervention assessment studies (e.g. (Roozenbeek et al., 2022; Spampatti et al., 2024)), including media literacy/misinformation games (Bradstreet and Starck, 2023; Wells, 2024). In the current study, we applied the MIST-20 version, which includes 20 items. Participants were asked to rate each presented item as either a ”fake” or ”real” news headline.
Literacy Assessment: New Media Literacy Scale. To assess changes in media literacy, we used the New Media Literacy Scale (NMLS), developed by Koc and Barut (Koc and Barut, 2016). The scale is designed to measure literacy in ”new media” (digital media and social networks) and is based on the four factors from the model by Chen et al.: functional consuming, critical consuming, functional prosuming, and critical prosuming (Chen et al., 2011). Our motivation for choosing the scale was due to the game being designed to actively engage users in content generation, and thus it is necessary to assess both the changes in the user’s information consumption and in their information prosumption (production and consumption). The NMLS is the only scale which provides this dimension, compared to other Media Literacy assessing instruments like ((Eristi and Erdem, 2017; Vraga et al., 2015; Ashley et al., 2013)) Additionally, the scale was developed and validated on the group of university students, which reflected our projected sample (mostly a young population with university degrees). The questionnaire includes 35 questions, rated on a 5-point Likert scale, ranging from ”strongly disagree” to ”strongly agree.”
@@
Behavioural Assessment: Verifying Online Information Scale (VOI - 7). To assess the effect of the game on the verification practices performed by the participants, we adapted the Verifying Online Information self-report scale (VOI) proposed by Tifferet (Tifferet, 2021). To the best of our knowledge, this is the only existing scale which focuses on the behaviors (verification practices) a person can adopt to verify the news. The questionnaire measures individuals’ differences in applying direct and indirect verification practices for online information, allowing us to track expected behavior changes in verification practices. We used the VOI-7 version, which demonstrated comparable construct characteristics to the original 22-question version while allowing to be completed more rapidly. The parameters were measured on a slider from 0 to 100, where participants were asked to indicate their likelihood of applying verification practices (Tifferet, 2021).
@@
Self-efficacy Assesement: Fake News Self-efficacy Scale. To measure perceived self-efficacy in dealing with fake news, we used a 3-item questionnaire developed by Hopp (Hopp, 2022). This questionnaire assessed participants’ confidence in three key areas: (1) their ability to identify news-like information that may be intentionally misleading, (2) their ability to distinguish between fake news and content produced with honest intentions, and (3) their ability to recognize news that may be unintentionally incorrect (i.e., misinformation). We chose the scale as a better alternative to the non-validated single-item measurement of confidence in identifying fake news, used by (Hinsley and Holton, 2021). Each item was rated on a seven-point scale, from ”strongly disagree” to ”strongly agree”.
4.1.2. Semi-structured Interview
@@
To evaluate the user’s experience in-depth, connected with the content of the game and the strategies implemented by users, we developed a protocol for a semi-structured interview. This protocol includes questions about the general experience, the perceived goal of the game, the perception of the opponent’s strategies, and the individual’s perception of the game’s effectiveness or ineffectiveness. The guidelines for the semi-structured interview are presented in the Supplementary material.
4.1.3. Log analysis
@@
The game logs collected the data including player-generated content, the time spent in each round, API responses showing the public opinion of different personas, trust level scores, and in-game events such as the amount of money players had, how much they spent, and what hints they purchased. This data provided a transcript of each session, enabling the research team to analyze players’ strategies, in-game behaviors, and decision-making processes.
4.1.4. Qualitative data analysis
@@
We employed a combined inductive-deductive approach to analyze the qualitative interview transcripts and gameplay logs(Kuckartz and Rädiker, 2019). This approach ensured a comprehensive understanding of the gameplay experience. Our primary objectives were to gain understanding of how participants perceived and understood misinformation through the game, how they learned to distinguish and apply debunking strategies during gameplay, and how interactions with other players influenced their behavior and learning. The analysis process began with inductive coding. Two researchers independently coded a subset of the data, identified themes, and then discussed and reconciled any coding discrepancies, iterating on the coding system as needed. Once the coding system was established, the two researchers independently coded the full dataset. A third researcher then reviewed the coded data, and any differences in interpretation were discussed until a consensus was reached.
@@ 4.2. Recruitment and Participants
Participants were recruited through flyers distributed in university-affiliated online media groups. We also encouraged participants to share information about the study within their social media networks using a snowballing technique. The eligibility criteria required participants to be adults and have sufficient English proficiency to play the game (we also do not forbid using translation engines if any of the aspects of the game are not understandable). Given that the proliferation of online misinformation is a global challenge and commonly reaches unsuspecting users(Ferrara et al., 2020), even people encountered misinformation, they are not necesserily knows about it if the information was not debunked later. Thus we did not require participants to have prior exposure.
60 participants initially expressed interest in the study by completing an online questionnaire that collected demographic information and availability. Ultimately, 47 participants were selected, forming 24 pairs for the game sessions. In one of the pairs, one of the study’s authors participated in a player role due to scheduling reasons. Because the data collection form allowed participants to skip questions they preferred not to answer, we noticed that 3 participants did not complete the entire MIST questionnaire, and 5 participants left some questions blank in the pre-procedural VOI questionnaire. Therefore, their data were excluded from the VOI and MIST data analyses. The study sample had the following characteristics: the participants’ ages ranged from 20 to 57, with a mean age of 25.87 years (SD = 6.265). 28 participants identified as female, 18 as male, and 1 preferred not to disclose their gender. 3 participants reported having an Associate degree, 29 a Bachelor’s degree, and 15 a Master’s degree; all participants reported having Eastern Asian origin (see participant demographic information in A.1).
@@ 4.3. Procedure
Once a person expressed interest and agreed to participate, they completed an initial questionnaire, providing demographic information and responding to the misinformation-related questionnaires. To prevent participants from intentionally biasing their responses, the questionnaire was administered 7-10 days before the gameplay experiment. After participants confirmed completion of the questionnaire, we scheduled the gameplay sessions. These experimental sessions were conducted either online via the VooV Meeting application or in person at a university meeting room. At the start of their test session, participants were given information sheets and consent forms to review and complete at their own pace. Once completed, Participants were introduced to the game setup and roles, and when they decided between themselves which role they would like to play. After the gaming session, participants again filled out the questionnaires. Finally, we conducted a short semi-structured interview to discuss their perceptions of the game. The entire session last approximately one and a half hours (See figure 7).
Figure 7. Overview of study procedure.
4.4. Ethical Consideration
The experimental design was approved by the Ethics Review Panel of (ANONYMIZED). As the game story was centered around a fictional pandemic, we informed participants about the theme in the consent form and asked them not to participate in the study if they perceived the topic of health/diseases to be disturbing. All participants gave their informed consent and were compensated 40 Chinese Renminbi upon completion of the game and interviews.
@@ 5. Results
5.1. Effect Game Play on Self Reported Discernment of Misinformation
@@ 5.1.1. Descriptive Statistics
Preliminary data analysis reveals significant violations of the normality assumption in pre-tested VOI and NMLS scales. Considering the rather small dataset, we decided to proceed with a non-parametric repeated measures approach (Related-Samples Wilcoxon Signed Rank Test to the results of the scales’ pre- and post-evaluations).
Effect of the educational game on the verification practicies (VOI-7) The test revealed significant differences between pre and post-gaming VOI scores (N = 42, Z = 4.361, p ¡ .001). The results suggested that the game positively affected the repertoire of used practices and/or the perceived will to use these practices.
@@
Effect of the educational game on Media Literacy To measure the effects of the game on Media Literacy, we first ran the Related-Samples Wilcoxon Signed Rank Test on the full scale. Then, to determine which components of Media Literacy were most affected by the game, we conducted separate subscale tests to analyze changes in each of the four subdomains of Media Literacy. The results demonstrated significant differences in Media Literacy scale results (N = 47, Z = 3.083, p = .002). The analysis revealed the following differences: the game significantly improved both functional consuming Z = 2.064, p = .039 and critical consuming Z = 3.344, p ¡.001 ), but not the functional prosuming Z = .435. p = .664 and critical prosuming Z = 1.868, p = .062. Therefore, the results suggest the game improves Media Literacy in the domains connected to understanding the content of the media and being able to critically evaluate the content of the media; however, it has not significantly improved the ability to produce media content which can be influential to others and convey author’s ideas (Koc and Barut, 2016).
@@
Effect of the educational game on self-efficacy towards misinformation We did not find significant differences in self-efficacy between pre and post-game measurements (Z = .743, p = .458).
@@
Effect of the educational game on the ability to recognise misinformation We took the naive approach to calculate the MIST score, taking it as the sum of the right answers on all 20 questions (Maertens et al., 2024). The results showed that participating in the game significantly improved the participant’s ability to discriminate between fake and real news (Z = 2.702, p = .007)
5.2. Qualitative Results
In this section, we present participants’ perceptions and understanding of misinformation, including how they learned to generate it and identify it through gameplay, as well as the strategies they developed for debunking misinformation.
5.2.1. Raise Awareness of Misinformation though gameplay
Participants reported an increased awareness of misinformation through both the confrontational mechanics of the game and the news narratives presented within it. The game enhanced players’ awareness of misinformation in two key ways.

### v2

@@ 3.4.1. Multiplayer Setup
Figure 5. Interaction System. Player 1 creates the online game room as the host and Player 2 joins. After Player 1 sends a request to GPT-4o API, Player 2 can view it. Both players in the online room receive the API response, which is processed and updated on their own game screens. Then Player 2 takes their turn to input information, following the same process as Player 1. At the end of the round, all in-game events (including player inputs, API responses, time, etc.) are logged locally.
@@ 4.1. Evaluation Methods
4.1.1. Key Concepts and Measurements
This paper focuses on the effectiveness of the game as a prebunking intervention in enhancing participants’ skills to protect themselves against misinformation. Following Lewandowsky and Van der Linden (Lewandowsky and Van Der Linden, 2021), we conceptualise prebunking interventions as follows: interventions which contribute to a person’s resilience to misinformation via raising critical media literacy skills (media knowledge and critical assessment of information), resulting in the will and ability to apply this knowledge to make decisions about information veracity. In contrast, debunking interventions can be understood as any intervention, which aims to correct the person’s opinion about information post hoc (after the person was exposed to and believed in) specific misinformation.
In this paper, we distinguish between the concept of “debunking” (the practice of challenging or disproving misinformation), and “debunking interventions” (social media platforms’ actions taken to perform debunking for user). We do not discuss “debunking interventions” (e.g., content labelling (Lewandowsky and Van Der Linden, 2021) or Community Notes (Chuai et al., 2024)); instead, we focus on ways to improve individuals’ debunking skills as a result of prebunking interventions.
Based on the definitions of prebunking interventions, it has become clear that we need to assess the educational effect of the intervention in relation to a) media literacy skills/knowledge gained in the intervention, b) behavioral intentions to apply that knowledge, and c) the practical ability to use this knowledge to recognize misinformation. Together, it would help to understand the effectiveness of the intervention and estimate the increase in debunking skills of the participants. This approach is similar to Tifferet (Tifferet, 2021) taxonomy for measuring susceptibility to misinformation, which can be viewed as the inverse of the resilience fostered by prebunking.
Tifferet categorizes measurements into three main groups: performance tasks (how well users can discriminate between fake and real pieces of news), general media literacy assessment (how much a person knows about different aspects of misinformation), and behavioral assessment (how much a person would like to use different strategies to verify online information) (Tifferet, 2021). As Tifferet argues, these three aspects are complementary to understanding user’s susceptibility to misinformation and, therefore, evaluation of improvement in each of them could provide a full picture of our game’s efficiency. In addition, as the game was designed to present complex scenarios, we decided to evaluate how confident participants were in their ability to recognise misinformation both before and after the intervention.
In the study, we used the validated questionnaires dedicated to measuring each of these aspects.
@@ 4.1.2. Questionnaires
Media Literacy Assessment: New Media Literacy Scale (NMLS). To assess changes in media literacy, we used the NMLS (Koc and Barut, 2016). The scale is designed to measure literacy in “new media” (digital media and social networks) and is based on the four factors from the model by Chen et al.: functional consuming, critical consuming, functional prosuming, and critical prosuming (Chen et al., 2011).
The NMLS was selected primarily because it assesses the critical dimensions of content production and consumption, which are core competencies our intervention aimed to establish. Notably, compared to other media literacy instruments (Eristi and Erdem, 2017; Vraga et al., 2015; Ashley et al., 2013), the NMLS is the only scale which provides this dimension. Additionally, the scale was developed and validated on a group of university students, which reflected our projected sample. The questionnaire includes 35 questions, rated on a 5-point Likert scale. We expected our intervention to enhance critical prosuming skills (as players were required to create influential content and sway LLM opinions) and critical consuming skills (as players had to analyze information from the game and other players to craft effective responses). We also anticipated positive effects on lower-level functional production and consumption skills (basic understanding of media consumption and creation), as the game offered basic training in reading, understanding, and responding comprehensively to media texts.
@@
Behavioural Assessment: Verifying Online Information Scale (VOI - 7). To assess the effect of the game on the verification practices performed by the participants, we adapted the VOI proposed by Tifferet (Tifferet, 2021). To the best of our knowledge, this is the only existing scale which focuses on the behaviors (verification practices) a person can adopt to verify the news. The questionnaire measures individuals’ differences in applying direct and indirect verification practices for online information, allowing us to track expected behavior changes in verification practices. We used the VOI-7 version, which demonstrated comparable construct characteristics to the original 22-question version while allowing to be completed more rapidly. The parameters were measured on a slider from 0 to 100, where participants were asked to indicate their likelihood of applying verification practices. As our game learning materials and the gaming procedure show the importance of verifying the information (via showing multiple misinformation-related events and presenting features of misinformation which should be checked to avoid being misled, we expected, that people will be more willing to apply verification practices.
Performance Assessment: Misinformation Susceptibility Test (MIST - 20). To assess changes in veracity discernment, we used the MIST (Maertens et al., 2024). To date, the MIST is the only fully validated misinformation susceptibility instrument. It takes into account the ability to recognize real news and fake news presented in equal proportion. The MIST framework is designed to allow for the comparison of results across different studies and interventions. The test has been implemented in multiple misinformation intervention assessment studies (e.g. (Roozenbeek et al., 2022; Spampatti et al., 2024)), including media literacy/misinformation games (Bradstreet and Starck, 2023; Wells, 2024) and having considerable predictive validity (Maertens et al., 2024), therefore giving comprehensive estimates of people’s ability to recognize real misinformation. We applied the MIST-20 version, which includes 20 items. Participants were asked to rate each item as either a “fake” or “real” news headline. We anticipated that participants would learn heuristics for identifying potential misinformation through intense interaction with the game. This interaction, which included hints about the characteristics of misinformation and the cognitive work of creating or debunking it, was expected to facilitate this learning.
@@
Self-efficacy Assesement: Fake News Self-efficacy Scale. To measure perceived self-efficacy in dealing with fake news, we used a 3-item questionnaire (Hopp, 2022). This questionnaire assessed participants’ confidence in three key areas: (1) their ability to identify news-like information that may be intentionally misleading, (2) their ability to distinguish between fake news and content produced with honest intentions, and (3) their ability to recognize news that may be unintentionally incorrect (i.e., misinformation). We chose the scale as a better alternative to the non-validated single-item measurement of confidence in identifying fake news, used by (Hinsley and Holton, 2021). Each item was rated on a seven-point scale.
4.1.3. Qualitative Data
@@
Semi-structured Interview: To evaluate the user’s experience in-depth, connected with the content of the game and the strategies implemented by users, we developed a protocol for a semi-structured interview. This protocol includes questions about the general experience, the perceived goal of the game, the perception of the opponent’s strategies, and the individual’s perception of the game’s effectiveness or ineffectiveness. The guidelines for the semi-structured interview are presented in the Supplementary material.
@@
Game Log: The game logs collected the data including player-generated content, the time spent in each round, API responses showing the public opinion of different personas, trust level scores, and in-game events such as the amount of money players had, how much they spent, and what hints they purchased. This data provided a transcript of each session, enabling the research team to analyze players’ strategies, in-game behaviors, and decision-making processes.
@@
Quantitative Data Analysis: We employed a combined inductive-deductive approach to analyze the interview transcripts and gameplay logs(Kuckartz and Rädiker, 2019). This approach ensured a comprehensive understanding of the gameplay experience. Our primary objectives were to gain understanding of how participants perceived and understood misinformation through the game, how they learned to distinguish and apply debunking strategies during gameplay, and how interactions with other players influenced their behavior and learning. The analysis process began with inductive coding. Two researchers independently coded a subset of the data, identified themes, and then discussed and reconciled any coding discrepancies, iterating on the coding system as needed. Once the coding system was established, the two researchers independently coded the full dataset. A third researcher then reviewed the coded data, and any differences in interpretation were discussed until a consensus was reached.
@@ 4.2. Recruitment and Participants
Participants were recruited through flyers and university-affiliated online media groups. We also encouraged participants to share information about the study within their networks. The eligibility criteria required participants to be adults and have sufficient English proficiency to play the game (we also do not forbid using translation engines if any of the aspects of the game are not understandable). Given that the proliferation of online misinformation is a global challenge and commonly reaches unsuspecting users(Ferrara et al., 2020). We did not require participants to have prior exposure.
60 participants initially expressed interest in participating in the game intervention study. Ultimately, 47 participants were selected, forming 24 pairs for the game sessions. In one of the pairs, one of the study’s authors participated in a player role due to scheduling reasons. Because the data collection form allowed participants to skip questions, 3 participants did not complete the entire MIST questionnaire, and 5 participants left some questions blank in the pre-procedural VOI questionnaire. Their data were excluded from the VOI and MIST data analyses. For the control condition, 57 participants signed up, and 50 successfully completed both the pre and post questionnaire and the reading exercise.
Demographic characteristics of the participants are summarized below. In the game intervention group, the participants’ ages ranged from 20 to 57, with a mean age of 25.87 years (SD = 6.265). 28 participants identified as female, 18 as male, and 1 preferred not to disclose their gender. 3 participants reported having an Associate degree, 29 a Bachelor’s degree, and 15 a Master’s degree; all participants reported having Eastern Asian origin. Participants in the control group ranged in age from 21 to 42, with a mean age of 27.46 (SD = 6.45). Thirty-three participants self-identified as female and 17 as male. Seventeen held a Master’s degree, 27 a Bachelor’s degree, and 6 an Associate degree. All participants reported being of East Asian origin. (See participant demographic information in Appendix A.1)
The experimental design was approved by the Ethics Review Panel of City University of Hong Kong. As the game story was centered around a fictional pandemic, we informed participants about the theme in the consent form and asked them not to participate in the study if they perceived the topic of health/diseases to be disturbing. All participants gave their informed consent and were compensated 40 Chinese Renminbi upon completion.
@@ 4.3. Procedure
Once a person expressed interest in participating, they completed an initial questionnaire containing demographic questions and measurement scales. To prevent participants from intentionally biasing their responses, the questionnaire was administered 7-10 days before the gameplay experiment. After participants confirmed completion of the questionnaire, we scheduled the gameplay sessions. These experimental sessions were conducted either online via the VooV Meeting application or in person at a university meeting room. At the start of their test session, participants were given information sheets and consent forms to review and complete at their own pace. Once completed, Participants were introduced to the game setup and roles, and when they decided between themselves which role they would like to play. After the gaming session, participants again filled out the questionnaires. Finally, we conducted a short semi-structured interview to discuss their perceptions of the game. The entire session last approximately one and a half hours (See Figure 6).
Figure 6. Overview of Game intervention study procedure.
4.4. Control Group
To verify whether the post-game improvements in the questionnaire results were not simply due to the Hawthorne effect (McCambridge et al., 2014), and to test whether passive exposure to misinformation alone could improve understanding, we conducted a control group study. Participants in the control group were asked to read examples of misinformation-containing social media content. The reading materials were selected from the Fake News Dataset(DBLP:conf/clef/KohlerSSWS0S22), which contains fact-checked claims and URLs from various fact-checking websites. We chose 40 posts related to the same topics used in the game intervention condition(themes regarding health, personal experience, alternative medicine, and pandemic). The estimated reading time was approximately one hour, which matched the expected duration of the game intervention. Before the task, we explicitly informed participants that the materials presented were examples of misinformation and that the purpose of reading them was to better understand how misinformation operates. Participants completed the same pre- and post-surveys used in the intervention group. The pre-survey was administered 7–10 days before the reading task. After completing the reading, participants took a short break, then proceeded to complete the post-survey.
@@ 5. Results
5.1. RQ1: Effect on the Ability to Recognise and Understand Misinformation
@@ 5.1.1. Descriptive Statistics
Preliminary data analysis reveals significant violations of the normality assumption in pre-tested NMLS scales. Considering the rather small dataset and the Likert-scales-based questionnaires used for most of the scales, we decided to proceed with a non-parametric repeated measures approach (Related-Samples Wilcoxon Signed Rank Test to the results of the scales’ pre- and post-evaluations). The scales’ descriptive statistics and normality tests results can be found in Table 2
Table 2. Descriptive Statistics and Normality Test Results for Variables Before and After Playing
Mean (SD) W DF Significance
Before After Before After Before After
Functional consuming 23.09 (0.528) 24.02 (0.466) 0.868 0.974 47 ¡.001 0.378
Critical consuming 44.04 (1.056) 47.74 (0.795) 0.863 0.971 47 ¡.001 0.285
Functional prosuming 28.02 (0.818) 28.64 (0.658) 0.898 0.954 47 ¡.001 0.060
Critical prosuming 36.04 (1.028) 37.91 (0.920) 0.941 0.958 47 0.019 0.092
Self-efficacy 15.17 (0.445) 15.26 (0.056) 0.970 0.959 47 0.265 0.101
VOI 416.79 (116.459) 485.93 (128.536) 0.978 0.960 42 0.588 0.153
MIST 11.5 (2.162) 12.5 (2.529) 0.969 0.965 44 0.284 0.208
@@
Effect of the Game on Media Literacy Skills: To measure the effects of the game on Media Literacy Skills, we first ran the Related-Samples Wilcoxon Signed Rank Test on the full scale. Then, to determine which components of Media Literacy were most affected by the game, we conducted separate subscale tests to analyze changes in each of the four subdomains of Media Literacy. The results demonstrated significant differences in Media Literacy scale results (N = 47, Z = 3.083, p = .002). The analysis revealed the following differences: the game significantly improved both functional consuming Z = 2.064, p = .039 and critical consuming Z = 3.344, p ¡.001 ), but not the functional prosuming Z = .435. p = .664 and critical prosuming Z = 1.868, p = .062. Therefore, the results suggest the game improves Media Literacy in the domains connected to understanding the content of the media and being able to critically evaluate the content of the media; however, it has not significantly improved the ability to produce media content which can be influential to others and support author’s ideas (Koc and Barut, 2016).
In contrast, we did not find the effect of intervention in the control group on any of the parameters media literacy: functional consuming (N = 50, Z = .991, p = .322), critical consuming (N = 50,Z = 1.505, p = .132), functional prosuming (N = 50,Z = 1.474, p = .141) and critical prosuming (N = 50, Z = 1.808, p =.072), showing that merely demonstrate the misinformation content is probably not enough to raise critical approach to media
Effect of the Game on the Verification Practices (VOI-7): The test revealed significant differences between pre and post-gaming VOI scores (N = 42, Z = 4.361, p ¡ .001). The results suggested that the game positively affected the repertoire of used practices and/or the perceived will to use these practices.
However, we also found significant differences between pre- and post-scores in the control group intervention (N = 50, Z = 2.208, p = .027), meaning exposure to the misinformation examples also makes people more vigilant and support checking intentions
@@
Effect of the Game on Self-efficacy towards Misinformation: We did not find significant differences in self-efficacy between pre and post-game measurements (N = 47, Z = .743, p = .458). In contrast,in control group, the intervention significantly self-efficacy (N=50, Z = 2.348, p = .019); in relation with the data of MIST we interpret these data as a tendency to be overconfident (more in Discussion section).
@@
Effect of the Game on the Ability to Recognise Misinformation: We took the “naive” approach to calculate the MIST score, taking it as the sum of the right answers on all 20 questions (Maertens et al., 2024). The results showed that participating in the game significantly improved the participants’ ability to discriminate between fake and real news (N = 44, Z = 2.702, p = .007) The participants in the control group did not demonstrate improvement in discriminating between fake and real information (N=50, Z = 1.277, p = .202 ).
5.1.2. Qualitative Results of Game Effects on Understanding Misinformation
Identifying Misinformation through Source Evaluation: After the gameplay sessions, participants reported increased awareness of the varying credibility of different information sources. The game helped them to realize that producers of misinformation often seeks to enhance credibility by deliberately referencing authoritative organizations. One participant reflected on this realization:
Identifying Misinformation through Emotional Manipulation Tactics: Participants learned various tactics for both creating and debunking misinformation through the game’s instructions and their in-game experiences. A particularly commonly identified tactic was emotional manipulation, which was noted by 35 out of 47 participants (18 Influencers and 17 Journalists). By analyzing the game logs, we identified common emotional manipulation strategies used in the game. Most players crafted messages designed to evoke anxiety and fear, while some also attempted to generate feelings of hope. For example, Influencer spread rumors about a doctor’s death, which directly incited public panic (N1). Across all rounds, Influencer frequently used emotional appeals and personal stories to enhance the perceived credibility of the misinformation (N6, N8). Additionally, invoking cultural pride and heritage was a powerful tactic used to build trust in misinformation (N4), while celebrity endorsements further increased the complexity and believability of the misinformation (N9). As illustrated in Figure 7, players significantly increased public trust in their information by using emotionally charged language (N31). These strategies align closely with the characteristics of misinformation, where emotional appeals are commonly used to influence public opinion (Chuai et al., 2022).
Interestingly, when playing the role of the Journalist, participant reflected on the emotionally inflammatory language used by the Influencers and helped them develop a more clear strategy for addressing misinformation. This approach involved separating the factual content of a message from its emotional manipulations and focusing more on the factual aspects, as one participants explained:
Critical Thinking about Misinformation Motives: After the gameplay sessions, many participants (31 out of 47) demonstrated an enhanced understanding of the intricacy of information and the varied perspectives it can convey. This experience increased their awareness of the importance of considering the motivations behind messages. Participants also mentioned that they are now more inclined to think critically about the goals behind the information they encounter, especially in real-life situations where such considerations are common. One participant explained how the game illustrated the pre-determined nature of many messages:
Impact of Gameplay on Future Debunking Actions: A few participants (4/47) shared that the game increased their likelihood of taking action against misinformation in the future. This change in attitude was driven either by participants’ previous negative personal experiences with misinformation or by their realization during gameplay of the serious consequences misinformation can have. The gameplay experience enhanced their willingness to invest time and effort into distinguishing and debunking false information. As one participant said:
5.2. RQ2: Player Behaviors in Response to Game Mechanics and Opponent Tactics